Skip to content

OSG Technology Area Meeting, 28 August 2017

Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl, Derek, Edgar, Suchandra, TimC

Announcements

Triage Duty

  • This week: Edgar
  • Next week: Derek
  • 9 (-3) open tickets

JIRA

# of tickets Δ State
154 -3 Open
20 -1 In Progress
14 +12 Ready for Testing
0 +0 Ready for Release

Release Schedule

Name Version Development Freeze Package Freeze Release Notes
September 3.4.3, 3.3.28 2017-08-28 2017-09-05 2017-09-12 5 week cycle
October 3.4.4, 3.3.29 2017-09-25 2017-10-02 2017-10-10
November 3.4.5, 3.3.30 2017-10-30 2017-11-06 2017-11-14 5 week cycle

Notes: Additional “urgent” releases may be scheduled for the 4th Tuesday of each month. The Testing date is when acceptance testing will be scheduled for releasable packages; if a package is added after this date, it may not be possible to schedule adequate testing time, thereby forcing it into the next release.

OSG Software Team

3.4.3/3.3.28

  • Tickets not marked RFT
Owner # tickets
Mat 10
Brian 7
Carl 5
Marian 1

Documentation

  • 6 documents migrated and meeting-related pages archived last week:
  • Derek has two students that are beginning to familiarize themselves with the migration process
  • We have enough experience with the process that we can start focusing on high-priority Release3 documents

Discussions

None this week

Support Update

  • Clemson (BrianL, Marian): expired WN CRLs and possibly an HTCondor IPv6 bug
  • JINR (BrianL): CE jobs held ("Error parsing classad or job not found") due to missing pbs_pro in /etc/blah.config
  • Issues with BLAHP at some of the BOSCO-CE's. Mats and Derek are continueing to investigate. Some issue with a sub-directory of the Bosco sandbox not being populated with the X509 certificate.

OSG Release Team

  • Suchandra Thapa is handling the September 12th release
  • Development Freeze today!
  • TimT may call upon software team members for testing assistance
3.3.28 Δ Both Δ 3.4.3 Δ Total Δ Status
2 -3 9 -2 0 -1 11 -6 Open
3 +1 8 +2 1 +0 12 +3 In Progress
5 +5 6 +5 3 +3 14 +13 Ready for Testing
0 +0 0 +0 0 +0 0 +0 Ready for Release
10 +3 22 +4 4 +2 37 +10 Total

Discussions

None this week

OSG Investigations Team

Last Week

  • GRACC-ITB moving forward
  • Some GRACC support of a slurm_meter issue, still debugging.
  • Indexing Glidein Logs in GRACC's ES - Ongoing
  • Backup Reports are now running every Monday (or, Sunday night, because GRACC node is in UTC time)
  • First step of creating ES snapshots for backup up GRACC - Ongoing.

This Week

  • GRACC backup of Dashboards
  • Initiate backups of ES snapshots
  • Start indexing GOC server status in GRACC ES
  • Work no Glidein logs in ES
  • Fix naming issues in GRACC related to explosion of "Fake" sites in records.

Ongoing

Discussions

XRootD developers have moved to a release model where they seek explicit sign-off from stakeholders